OpenStudy (anonymous):

Find the slope of the line through P and Q. P(4, 3), Q(−1, 0)

OpenStudy (anonymous):

y1-y=m(x1-x) m=(3-0)/(4-(-1)) m=...?

OpenStudy (amistre64):

its the ratio of how far y moves with respect to how far x moves

OpenStudy (amistre64):

y moves from 3 to 0; -3 x moves from 4 to -1; -5 slope is the ratio of -3 to -5

OpenStudy (anonymous):

so what is the slope lol i dont understand what you are saying

OpenStudy (amistre64):

a ratio is another way to say "its a fraction".

OpenStudy (anonymous):

so -3/5

OpenStudy (anonymous):

?

OpenStudy (amistre64):

close :) the negatives cancel each other out and your left with 3/5

OpenStudy (anonymous):

gotcha so the answer is 3/5

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

what about for this one P(3, −9), Q(−3, 8)

OpenStudy (amistre64):

If we view these as P(x,y) and Q(x,y) the ratio is then: (Py-Qy)/(Px-Qx)

OpenStudy (chaise):

Subtract any x value from the other x value. Subtract the y value from the other y value. Divide y by x, there is your gradient, or slope. You can further write an equation for this line by substituting in your x and y values from any point which the graph goes through, to find your "b" in the equation "y = mx +b"

OpenStudy (anonymous):

-72/9?

OpenStudy (amistre64):

another way is simply to subtract one point from the other ( 3,-9) -(-3, 8) -------- which is the same thing as this: (3,-9) +(3,-8) -------- 6,-17 ; now stack the y part over x part

OpenStudy (anonymous):

so -17/6

OpenStudy (amistre64):

yep; and they like it to be simplified .... but i think thats as simplified as it gets

OpenStudy (anonymous):

thanks alot